OpenStudy (andrewthecookie):

There are 1296 in (squared) in one square yard. Use a proportion to find out how many square yards of carpet it would take to over the floor

OpenStudy (andrewthecookie):

1/4 = 18 inches (scale)

OpenStudy (andrewthecookie):

@CausticSyndicalist

OpenStudy (andrewthecookie):

Halp xc

OpenStudy (andrewthecookie):

@Compassionate

OpenStudy (andrewthecookie):

Heres a photo if it helps

OpenStudy (andrewthecookie):

@TheSmartOne

OpenStudy (andrewthecookie):

Sorry but hurry :I

OpenStudy (andrewthecookie):

@zepdrix What are the acutal dimensions of the sofa in inches (look at picture)

OpenStudy (andrewthecookie):

@Hero

OpenStudy (wolf1728):

That drawing does NOT have a scale that is accurately stated. It would be better if they said each space between the lines is 18 inches. That chart is 12 lines by 10 lines which means the room is 12 * 18 inches = 216 inches (18 feet) by 10 * 18 inches = 180 inches (15 feet) So, it is 18*15 = 270 square feet = 30 square yards I didn't set up any proportions but you have the answer.

OpenStudy (andrewthecookie):

:O Ty!

OpenStudy (wolf1728):

Okay :-)

OpenStudy (andrewthecookie):

Answer the second plz?

OpenStudy (andrewthecookie):

the chioces are 90 x 45 and 90x 36 and 75 x 25 and 80 x 30

OpenStudy (andrewthecookie):

srry for rushing but just tell me because i have to get this done

OpenStudy (wolf1728):

Well the room is 216 by 180 inches and that does not seem to be the answer.

OpenStudy (andrewthecookie):

which one is it?

OpenStudy (wolf1728):

I have no idea - those products (90 x 45) aren't in the same proportion and none of those = 38,880 square inches